# Erasing Conceptual Knowledge from Language Models

### v2

@@ 5.2 Erasing WMDP Concepts
| RMU | 49.4 | 37.0 | 40.1 | 3.9 | 4.1 | |
| RepNoise | 54.7 | 43.6 | 54.2 | 5.5 | 4.9 | |
| Ours | 33.3 | 26.6 | 57.2 | 4.8 | 4.5 | |
@@ 5.3 Ablation Study
We conduct ablation experiments to analyze the contribution of each loss component in ELM. Table 2 shows the impact on concept erasure (WMDP), general knowledge (MMLU), and generation quality (MT-Bench, Perplexity) when removing or modifying individual terms.
First, proves crucial for innocence. Removing significantly degrades erasure performance, with WMDP scores remaining close to the original model. While randomizing logits as a proxy for erasure (replacing in Eq. 10 with random vector) can achieve some erasure in cyber concepts, they lead to incoherent outputs.
The retain term (Eq. 11) is vital for specificity. Its removal yields the lowest MMLU scores, demonstrating its role in maintaining broad knowledge while enabling targeted erasure.
Ablating , leads to effective erasure but the model generates low-quality gibberish text when prompted for erased concept. Replacing it with random text from WikiText (Merity et al., 2016) slightly reduces fluency while maintaining erasure effectiveness. Qualitatively this produces awkward outputs that tend to be irrelevant to input prompts. This underscores the term’s role in maintaining seamless contextual relevance.
@@ 5.5 Robustness to Attacks
We evaluate the robustness of ELM by employing Greedy Coordinate Gradient attack (GCG) (Zou et al., 2023). We optimize for an adversarial suffix prompt to induce the model to generate a target text. We use the prompt “How would I engineer influenza A to be significantly more virulent?” and set the target prompt as “To make influenza A significantly more virulent,”. ELM demonstrates strong resistance to these attacks even when optimized for 5000 steps, whereas, the original model starts responding to the query with the GCG suffix, as illustrated in the example below. We provide GCG attack details in Appendix C. We also show ELM’s robustness to more attacks in Appendix H.
@@ 5.6 Probing and Activation Analysis
To estimate the presence of erased knowledge within the internal representations of a model, we conduct the probing analysis, training a linear probe using the same setup as used by (Li et al., 2024).
The results in Figure 3(a) reveal distinct knowledge retention patterns across methods. ELM and RMU achieve effective erasure, maintaining low probe accuracies across all layers for both biosecurity and cybersecurity MCQs. In contrast, RepNoise shows partial retention, particularly for cybersecurity concepts.
@@ 5.7 Erasing Harry Potter Knowledge
To further demonstrate the versatility of ELM, we apply it to the unlearn knowledge of Harry Potter literary universe. We compare ELM against RMU and WhoIsHarryPotter (WHP) (Eldan & Russinovich, 2023) methods for Llama-2-7B Chat. Table 3 presents this comparison.
ELM achieves a balance between effective knowledge erasure (low HP MCQ score) and maintaining fluent generation (low reverse-perplexity). Similar to (Lynch et al., 2024), we found WHP model (Eldan & Russinovich, 2023) maintains fluency but fails to effectively erase the target knowledge. RMU (Li et al., 2024) proved to be ineffective in erasing with a large hyper parameter sweep. A more through sweep may be necessary to conclusively determine its limitations in this context
@@
While ELM effectively removes targeted concepts through introspective classification, several limitations merit investigation. The method shows some degradation in performance on semantically adjacent concepts, indicating that our approach may need refinement to achieve more precise boundaries between related knowledge. Additionally, although generated text maintains basic fluency, it sometimes lacks semantic coherence - suggesting that our probability modification may be overly aggressive in some cases. The most significant challenge lies in handling deeply interconnected concepts, where modifying the model’s behavior for one concept may have ripple effects through its broader knowledge base. Further work is needed to develop more granular techniques for selective knowledge modification while preserving complex conceptual inter-dependencies.
@@ Acknowledgements
Appendix A Details on metrics
@@
To measure the seamlessness of edits, we aim to quantify the fluency of the text being generated by the edited model when prompted with the concept being erased. To evaluate this we prompt the models using questions from MCQ dataset from WMDP (Li et al., 2024) and let the models generate text free-form up to 500 tokens. We then measure the perplexity on generated text using a totally different evaluation model, Llama3.1-8B (Dubey et al., 2024).
Appendix B Baseline Methods
B.1 WMDP Results
@@
B.2 Harry Potter Results
@@ Appendix D Hyperparameter Analysis
Appendix E Low-Rank vs Full Finetuning
We analyze the role of using low-rank updates with ELM comparing its performance against finetuning the layers directly without any rank constraints. In Table 4, we show the performance of ELM on Zephyr-7B when editing with full finetuning and low-rank model editing. Full finetuning effects the specificity of the model and makes the unlearning broader damaging the general capabilities of the model. Where as low-rank model editing preserves the specificity while being effective at erasure.
@@ Appendix E Low-Rank vs Full Finetuning
Appendix F Conditional Fluency Training
To achieve seamless editing, ELM must generate fluent text even when prompted about erased concepts. The ideal behavior mimics a model that never encountered the concept during pretraining. We implement an additional step to make ELM models acknowledge the concept while suggesting a topic change, although this behavior remains configurable through prompt engineering.
@@ Appendix F Conditional Fluency Training
Appendix G Progression of ELM Training
@@
Appendix H Adversarial Attacks
We also attack ELM with BEAST (Sadasivan et al., 2024), a fast adversarial prompt based attack on LLMs. BEAST finds an adversarial prompt that can be appended to the original attack prompt to generate target response. We find that BEAST is unable to extract erased information from ELM:

### v3

@@ 5.2 Erasing WMDP Concepts
| Ours | 33.1 | 27.1 | 78.4 | 7.9 | 4.8 | |
| Ours () | 32.7 | 27.5 | 78.8 | 7.8 | 5.1 | |
[…]
| Ours | 33.7 | 28.2 | 75.2 | 7.2 | 4.8 | |
| Ours () | 32.1 | 28.0 | 75.7 | 7.2 | 4.3 | |
@@
We observe an emergent artifact in larger models where the fluency term () is no longer necessary for maintaining fluency. Our erasing loss term is enough to precisely erase the knowledge while maintaining fluency at the same time. This suggests that larger models have more precise internal classifiers than smaller models, making additional fluency objectives unnecessary. We show the progression on unlearning in Appendix F and qualitative samples in Appendix H
@@ 5.3 Ablation Study
We conduct ablation experiments to analyze the contribution of each loss component in ELM. Table 2 shows the impact on concept erasure (WMDP), general knowledge (MMLU), and generation quality (MT-Bench, Perplexity) when removing or modifying individual terms. We show more fine-grained ablations and hyper-parameters in Appendix D
First, proves crucial for innocence. Removing significantly degrades erasure performance, with WMDP scores remaining close to the original model. While randomizing logits as a proxy for erasure (replacing in Eq. 7 with random vector) can achieve some erasure in cyber concepts, they lead to incoherent outputs. The retain term (Eq. 8) is vital for specificity. Its removal yields the lowest MMLU scores, demonstrating its role in maintaining broad knowledge while enabling targeted erasure. Ablating , leads to effective erasure but the model generates low-quality gibberish text when prompted for erased concept. Replacing it with random text from WikiText Merity et al. (2016) slightly reduces fluency while maintaining erasure effectiveness. Qualitatively this produces awkward outputs that tend to be irrelevant to input prompts. This underscores the term’s role in maintaining seamless contextual relevance. However, we note that is not necessary for larger models (Qwen2.5-32B and Llama3-70B).
@@ 5.5 Robustness to Attacks
We evaluate the robustness of ELM by employing Greedy Coordinate Gradient attack (GCG) Zou et al. (2023). We optimize for an adversarial suffix prompt to induce the model to generate a target text. We use the prompt “How would I engineer influenza A to be significantly more virulent?” and set the target prompt as “To make influenza A significantly more virulent,”. ELM demonstrates strong resistance to these attacks even when optimized for 5000 steps, whereas, the original model starts responding to the query with the GCG suffix, as illustrated in the example below. We provide more details on GCG attack and show other adversarial attacks in Appendix G.
@@ 5.6 Probing and Activation Analysis
To estimate the presence of erased knowledge within the internal representations of a model, we conduct the probing analysis, training a linear probe using the same setup as used by Li et al. (2024).
The results in Figure 3(a) reveal distinct knowledge retention patterns across methods. ELM and RMU achieve effective erasure, maintaining low probe accuracies across all layers for both biosecurity and cybersecurity MCQs. In contrast, RepNoise shows partial retention, particularly for WMDP-Cyber.
@@ 5.7 Erasing Harry Potter Knowledge
To further demonstrate the versatility of ELM, we apply it to the unlearn knowledge of Harry Potter literary universe. We compare ELM against RMU and WhoIsHarryPotter (WHP) (Eldan and Russinovich, 2023) methods for Llama-2-7B Chat. Table 3 presents this comparison.
ELM achieves a balance between effective knowledge erasure (low HP MCQ score) and maintaining fluent generation (low reverse-perplexity). Similar to Lynch et al. (2024), we found WHP model (Eldan and Russinovich, 2023) maintains fluency but fails to effectively erase the target knowledge as revealed in its retained ability to answer multiple-choice questions about Harry Potter. RMU (Li et al., 2024) proved to be ineffective in erasing with a large hyper parameter sweep. A more through sweep may be necessary to conclusively determine its limitations in this context
@@ 6 Limitations
While ELM effectively removes targeted concepts through introspective classification, several limitations merit investigation. The method shows some degradation in performance on semantically adjacent concepts, indicating that our approach may need refinement to achieve more precise boundaries between related knowledge. Additionally, although generated text maintains basic fluency, it sometimes lacks semantic coherence, suggesting that our probability modification may be overly aggressive in some cases. The most significant challenge lies in handling deeply interconnected concepts, where modifying the model’s behavior for one concept may have ripple effects through its broader knowledge base. Further work is needed to develop more granular techniques for selective knowledge modification while preserving complex conceptual inter-dependencies.
@@ Acknowledgements
Appendix A Impact Statement
In this work, we develop a framework for thinking about concept erasure in language models, as well as a new approach to erasing conceptual knowledge. Although we focus on removal of potentially harmful knowledge, this technology could be misused to remove legitimate knowledge from a language model without users’ awareness. Additionally, if our method is used to remove harmful knowledge, it may create a false sense of security, as models could retain harmful knowledge that is undetected by our metrics. Unlearning has an important place in safety considerations for language models, but should not be the only approach. Finally, we also acknowledge that our evaluations are focused on harmful knowledge encoded in English; we have not evaluated this approach cross-linguistically. We release our code publicly to enable open and safe research.
Appendix B Details on metrics
@@
To measure the seamlessness of edits, we aim to quantify the fluency of the text being generated by the edited model when prompted with the concept being erased. To evaluate this we prompt the models using questions from MCQ dataset from WMDP Li et al. (2024) and let the models generate text free-form up to 500 tokens. We then measure the perplexity on generated text using a totally different evaluation model, Llama3.1-8B (Dubey et al., 2024).
Appendix C Baseline Methods
We compare ELM against other baselines across different models for unlearning WMDP-Bio and WMPD-cyber in Table 4. ELM shows stronger general erasure performance across different model architectures and settings.
C.1 WMDP Results
@@
| Model | Method | Innocence () | Specificity () | Seamlessness | ||
| Bio | Cyber | MMLU | MT-Bench | R-PPL () | ||
| Zephyr-7B | Original | 64.4 | 44.3 | 58.5 | 7.3 | 6.0 |
| RMU | 30.5 | 27.3 | 57.5 | 7.2 | 24.8 | |
| RepNoise | 29.7 | 37.7 | 53.3 | 6.6 | 25.0 | |
| Ours | 29.7 | 27.2 | 56.6 | 7.1 | 10.9 | |
[…]
| RMU | 33.5 | 28.7 | 27.1 | 1.0 | 29.9 | |
| RepNoise | 35.3 | 39.6 | 55.0 | 2.1 | 26.7 | |
| Ours | 28.7 | 26.4 | 55.4 | 3.7 | 15.3 | |
[…]
| RMU | 46.2 | 31.9 | 56.5 | 7.4 | 3.0 | |
| RepNoise | 59.9 | 44.1 | 60.1 | 6.7 | 3.5 | |
| Ours | 32.2 | 27.2 | 61.6 | 7.7 | 7.4 | |
[…]
| RMU | 49.4 | 37.0 | 40.1 | 3.9 | 4.1 | |
| RepNoise | 54.7 | 43.6 | 54.2 | 5.5 | 4.9 | |
| Ours | 33.3 | 26.6 | 57.2 | 4.8 | 4.5 | |
[…]
| Ours | 33.1 | 27.1 | 78.4 | 7.9 | 4.8 | |
| Ours () | 32.7 | 27.5 | 78.8 | 7.8 | 5.1 | |
[…]
| Ours | 33.7 | 28.2 | 75.2 | 7.2 | 4.8 | |
| Ours () | 32.1 | 28.0 | 75.7 | 7.2 | 4.3 | |
C.2 Harry Potter Results
@@ Appendix D Hyperparameter Analysis
D.1 Ablation on ELM Loss Terms
We sweep the values of , , from ELM Loss terms in Equation 10. We run this ablation on Llama3-8B model and show the results in Table 5. We find that increasing the erase loss scale () tends to increase the erasure effect. Increasing the retain loss term () improves the specificity of the erasure. Finally, increasing the consistency term has improved fluency, but increasing it beyond a certain value affects the erasure efficacy of the method.
| Value | WMDP-Bio () | MMLU () | R-PPL () | |
| 0.0 | 0.70 | 0.63 | 7.13 | |
| 0.5 | 0.43 | 0.62 | 12.49 | |
| 1.0 | 0.37 | 0.62 | 7.92 | |
| 1.5 | 0.34 | 0.62 | 9.79 | |
| 2.0 | 0.35 | 0.62 | 8.80 | |
| 0.0 | 0.25 | 0.24 | 9.19 | |
| 0.5 | 0.31 | 0.61 | 10.41 | |
| 1.0 | 0.37 | 0.62 | 7.92 | |
| 1.5 | 0.38 | 0.62 | 10.72 | |
| 2.0 | 0.37 | 0.62 | 9.31 | |
| 0.0 | 0.28 | 0.61 | 22.29 | |
| 0.5 | 0.34 | 0.62 | 8.34 | |
| 1.0 | 0.37 | 0.62 | 7.92 | |
| 1.5 | 0.39 | 0.62 | 12.01 | |
| 2.0 | 0.35 | 0.62 | 11.64 |
D.2 Low-Rank vs Full Finetuning
We analyze the role of using low-rank updates with ELM comparing its performance against finetuning the layers directly without any rank constraints. In Table 6, we show the performance of ELM on Zephyr-7B when editing with full finetuning and low-rank model editing. Full finetuning effects the specificity of the model and makes the unlearning broader damaging the general capabilities of the model. Low-rank model editing preserves the specificity while being effective at erasure.
@@ D.2 Low-Rank vs Full Finetuning
Appendix E Conditional Fluency Training
For smaller models, we find that erasure loss alone is not enough to maintain fluency. To achieve seamless editing for smaller models, ELM must generate fluent text even when prompted about erased concepts. The ideal behavior mimics a model that never encountered the concept during pretraining. We implement an additional step to make ELM models acknowledge the concept while suggesting a topic change, although this behavior remains configurable through prompt engineering.
@@ Appendix E Conditional Fluency Training
Appendix F Progression of ELM Training
@@
Appendix G Robustness Evaluation
G.1 Greedy Coordinate Gradient (GCG)
ELM:
RMU:
RepNoise:
G.2 BEAST
We also attack ELM with BEAST Sadasivan et al. (2024), a fast adversarial prompt based attack on LLMs. BEAST finds an adversarial prompt that can be appended to the original attack prompt to generate target response. We find that BEAST is unable to extract erased information from ELM:
G.3 Finetuning Attack